OpenStudy (anonymous):

if g(u)=(1)/(9u+6(1/2)) find g'(u)

OpenStudy (anonymous):

do i flip it first?

OpenStudy (lalaly):

\[g(u)=(9u+3)^{-1}\]

OpenStudy (anonymous):

why 3?

OpenStudy (lalaly):

is this 6 multiplied by 1/2?

OpenStudy (anonymous):

the 1/2 meant that everything in the denominator is under a radical

OpenStudy (lalaly):

ohh\[g(u)=\frac{1}{(9u+6)^{\frac{1}{2}}}\]?

OpenStudy (anonymous):

yeah...

OpenStudy (lalaly):

\[g(u)=(9u+6)^{-\frac{1}{2}}\]

OpenStudy (lalaly):

do u know how to differentiate this?

OpenStudy (anonymous):

multiply the first # by the power and reduce the power by 1

OpenStudy (lalaly):

\[g'(u)=-\frac{1}{2}(9u+6)^{-\frac{3}{2}}\times 9\]

OpenStudy (lalaly):

yeah then multiply it all by the derivative of whats inside the bracket

OpenStudy (anonymous):

lalaly knows it all :D

OpenStudy (anonymous):

wheres the 9 from?

OpenStudy (lalaly):

got it?

OpenStudy (lalaly):

lol fool no i dont

OpenStudy (anonymous):

Chain Rule

OpenStudy (anonymous):

yesss you doo! :D

OpenStudy (lalaly):

haha

OpenStudy (anonymous):

could someone explain the chain rule then? lol